# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ 2.2. Polymorphisms are injective
Define a relation $I_{4}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{(x,y,u,v)\mid x=y\Rightarrow u=v\}$ on a countably infinite set $V$. Then $\mathrm{CSP}(V;I_{4})$ is solvable in Datalog [24], and in particular does not pp-construct $\mathfrak{K}_{3}$; therefore, the polymorphisms of this template satisfy the pseudo-Siggers identity by the above-mentioned theorem of Barto and Pinsker. However, any polymorphism of $I_{4}$ must be injective up to dummy variables [17], and hence cannot satisfy the Siggers identity.
We say that a template is Pol-injective if all of its polymorphisms are essentially injective, i.e. injective up to dummy variables – this is the case if and only if the relation $I_{4}$ is invariant under its polymorphisms. As it turns out, the relation $I_{4}$ can be added to the templates constructed in Theorem 2.2 at no computational cost. It thus follows that any algebraic condition on polymorphisms capturing a complexity class closed under Datalog-reductions for CSPs of structures within the scope of Conjecture 1.4 must necessarily be satisfiable by essentially injective operations. Until now, this fact was understood by experts on an intuitive level, but our result provides a rigorous confirmation.
@@
The expansions $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ and $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$ of $\mathbb{Q}^{\mathfrak{A}}$ and $(\mathbb{Q};<)^{\mathfrak{B}}$, respectively, by the relation $I_{4}$ enjoy all properties of Theorem 2.2, except that CSP-injectivity is replaced by Pol-injectivity.
For finite structures the only essentially injective polymorphisms are constants and projections. It is therefore of no surprise that so few universal-algebraic conditions for CSPs from the finite have been successfully lifted to the infinite. One prominent example which can actually be lifted is the satisfiability of quasi near-unanimity identities, which captures bounded strict width both in the finite and in the infinite [18]; bounded strict width corresponds to solvability in a particular proper fragment of Datalog that is not closed under the Datalog-reductions in Theorem 2.2. Examples of conditions for polymorphism clones that can be satisfied by essentially injective operations are any pseudo height-1 identities (e.g., the pseudo-Siggers identity), but also the dissected near-unanimity identities [48] which were proposed in [34] as a possible candidate for solvability of CSPs in fixed-point logic.
Marimon and Pinsker have recently provided a negative answer, within the scope of Conjecture 1.4, to a question of Bodirsky [16, Question 14.2.6(27)] asking whether every $\omega$-categorical CSP template without algebraicity which is solvable in Datalog has a binary injective polymorphism [57, Corollary 8.10]. We contrast this negative result with a corollary to Theorem 2.4 that implies that up to Datalog reductions the answer is positive.
@@
Let $\mathfrak{A}$ be a non-trivial reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. If $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, then $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ (from Theorem 2.4) has a binary injective polymorphism.
2.3. Sandwiches are relevant
@@
One of the currently most vibrant branches of research in constraint satisfaction are Promise Constraint Satisfaction Problems (PCSPs). For two relational structures $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ such that $\mathfrak{S}_{1}$ maps homomorphically into $\mathfrak{S}_{2}$, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to decide whether a given finite structure $\mathfrak{J}$ in the same signature as $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ maps homomorphically into $\mathfrak{S}_{1}$ or does not even map homomorphically into $\mathfrak{S}_{2}$; the instance $\mathfrak{J}$ is promised to satisfy one of those two cases. Thus, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to compare $\mathrm{CSP}(\mathfrak{S}_{1})$ to a relaxation $\mathrm{CSP}(\mathfrak{S}_{2})$ of it, and since $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{1})=\mathrm{CSP}(\mathfrak{S}_{1})$, promise constraint satisfaction problems generalize the CSP framework. This generalization is vast, and contains many well-known computational problems such as APPROXIMATE GRAPH COLORING [46], $(2+\epsilon)$-SAT [2], and HYPERGRAPH COLORING [41]. Although formally not necessary, research has focused on finite templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$: this is justified for example by the fact that even for structures over a Boolean domain, there is no complete complexity classification yet. Hence, the two here considered extensions of finite-domain CSPs to $\omega$-categorical templates, or alternatively to PCSPs, are somewhat orthogonal, and it is natural to wonder whether there exist any connections, in particular of algorithmic nature.
@@
One potential link is the use of $\omega$-categorical CSP templates as sandwiches to prove tractability of PCSPs, as follows. Note that if $\mathfrak{A}$ is a structure such that $\mathfrak{S}_{1}$ maps homomorphically to $\mathfrak{A}$ and $\mathfrak{A}$ maps homomorphically to $\mathfrak{S}_{2}$, then $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ trivially reduces to $\mathrm{CSP}(\mathfrak{A})$; we call $\mathfrak{A}$ a sandwich for the template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$. This simple observation is more powerful than it might first appear: in fact, so far, every known tractable PCSP can be reduced to the tractable CSP of some sandwich structure [58, 40]. On the other hand, Barto [4] provided an example of a (finite-domain) PCSP template with an infinite polynomial-time tractable sandwich structure that is not finitely tractable, i.e. such that every finite-domain sandwich structure has an NP-complete CSP [4]. Since his sandwich structure is not $\omega$-categorical, he raised the question whether such an example could also be constructed with an $\omega$-categorical structure ([4, Question IV.1]). Subsequently the question arose, and was promoted in particular by Zhuk at the CSP World Congress 2023, whether structures within the scope of Conjecture 1.4 could ever serve as tractable sandwich structures in the absence of a finite one.
@@
In the present paper, we provide a strongly affirmative answer to this question and create non-finitely tractable finite-domain PCSPs from virtually every structure $\mathfrak{A}$ within the scope of the Bodirsky-Pinsker conjecture. Our construction consists of two steps. First, we show such construction is possible under the additional assumption that $\mathfrak{A}$ is equipped with an inequality predicate and that it is a reduct of a structure $\mathfrak{B}$ which is linearly ordered. We then combine this construction with Theorem 2.2. Modulo an extra step consisting of taking generic superpositions of structures with $(\mathbb{Q};<)$ (for which we need the property of no algebraicity), this finishes our proof.
Theorem 2.6.
@@
For every non-trivial reduct $\mathfrak{A}$ of a finitely bounded homogeneous structure $\mathfrak{B}$, there exists a reduct $\mathfrak{A}^{\prime}$ of a finitely bounded homogeneous structure $\mathfrak{B}^{\prime}$ and a finite PCSP template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that
@@
    $\mathrm{CSP}(\mathfrak{A})$ and $\mathrm{CSP}(\mathfrak{A}^{\prime})$ are Datalog-interreducible.
  3. (3)
    $\mathfrak{A}^{\prime}$ is a sandwich for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$.
  4. (4)
    Every finite sandwich for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ pp-constructs $\mathfrak{K}_{3}$
    (and hence $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is not finitely tractable).
@@
2.4. Related work
By a theorem of Kechris, Pestov, and Todorčević [54], a homogeneous structure $\mathfrak{B}$ with a countable relational signature is Ramsey if and only if its automorphism group $\mathrm{Aut}(\mathfrak{B})$ is extremely amenable, and hence the Ramsey property only depends on $\mathrm{Aut}(\mathfrak{B})$ viewed as an abstract topological group. They also observed that extreme amenability is preserved under taking wreath products; although they do not state this result explicitly, they do provide an example as a proof of concept: the automorphism group of the random convexly ordered equivalence relation. Scow [66, Theorem 5.13] does state this result, and essentially provides a full proof, modulo Theorem 4.7 in [54]. We kindly borrow her statement and in return provide a positive answer to one of her open questions regarding semi-retractions [12, Question 4.3]. In simple words, a semi-retraction is a certain purely algebraic construction allowing to transfer the Ramsey property from one structure to another.
Proposition 2.7.
The random linearly ordered graph is a semi-retract of the linearly ordered atomless Boolean algebra.
@@ 2.4. Related work
The first two examples of a tractable $\omega$-categorical sandwich for a finite-domain PCSP that is not finitely tractable have recently been provided in the preprint [59, Propositions 35 and 36]. Our results compare to the results of that work as follows: [59, Theorem 1] shows that every CSP within the scope of the Bodirsky-Pinsker conjecture is polynomial-time equivalent to the PCSP of a half-infinite template, more precisely a template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that $\mathfrak{S}_{1}$ is in the scope of the Bodirsky-Pinsker conjecture and $\mathfrak{S}_{2}$ is finite. In contrast, our Theorem 2.6 provides a non-finitely tractable finite-domain PCSP template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ sandwiching a given structure $\mathfrak{A}$ from Conjecture 1.4 up to Datalog interreducibility. Since our PCSP template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is finite and not half-infinite, we cannot guarantee polynomial-time equivalence between $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ and $\mathrm{CSP}(\mathfrak{A})$.
2.5. Organization of this article
We define the basic notions not defined in the introduction in Section 3. Then Section 4 will be devoted to Theorems 2.2 and 2.4; it also contains the definition of Datalog reductions. Section 5 contains a proof sketch of Theorem 2.6. Finally, we draw conclusions in Section 6. Due to space restrictions, we omit most technical parts of proofs or only sketch them; full proofs can be found in the appendix.
3. Preliminaries
@@
The set $\{1,\dots,n\}$ is denoted by $[n]$, and we use the bar notation $\bar{t}$ for tuples. For a function $f\colon A^{n}\rightarrow B$, the component-wise action of $f$ on $k$-tuples $(x_{1,1},\dots,x_{1,k}),$ $\dots$,… , $(x_{n,1},\dots,x_{n,k})\in A^{k}$ is defined as the $k$-tuple $f\big{(}(x_{1,1},\dots,x_{1,k}),\dots,(x_{n,1},\dots,x_{n,k})\big{)}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\big{(}f(x_{1,1},\dots,x_{n,1}),\dots,f(x_{1,k},\dots,x_{n,k})\big{)}$.
Relational structures.
@@
A (relational) signature $\tau$ is a set of relation symbols, each $R\in\tau$ with an associated natural number called arity. A (relational) $\tau$-structure $\mathfrak{A}$ consists of a set $A$ (the domain) together with the relations $R^{\mathfrak{A}}\subseteq A^{k}$ for each $R\in\tau$ with arity $k$. An expansion of $\mathfrak{A}$ is a $\sigma$-structure $\mathfrak{B}$ with $A=B$ such that $\tau\subseteq\sigma$ and $R^{\mathfrak{B}}=R^{\mathfrak{A}}$ for each relation symbol $R\in\tau$. Conversely, we call $\mathfrak{A}$ a reduct of $\mathfrak{B}$ and denote it by $\mathfrak{B}|^{\tau}$. For a positive integer $d$ and a $\tau$-structure $\mathfrak{A}$, the $d$-th power of $\mathfrak{A}$ is the $\tau$-structure $\mathfrak{A}^{d}$ with domain $A^{d}$ and relations $R^{\mathfrak{A}^{d}}=\{\big{(}(a_{1,1},\dots,a_{1,d}),\dots,(a_{k,1},\dots,a_{k,d})\big{)}\in(A^{d})^{k}\mid(a_{1,1},\dots,a_{k,1}),\dots,(a_{1,d},\dots,a_{k,d})\in R^{\mathfrak{A}}\}$ for each $R\in\tau$ of arity $k$. The substructure of a $\tau$-structure $\mathfrak{A}$ on a subset $B\subseteq A$ is the $\tau$-structure $\mathfrak{B}$ with domain $B$ and relations $R^{\mathfrak{B}}=R^{\mathfrak{A}}\cap B^{k}$ for every $R\in\tau$ of arity $k$. The factor of a $\tau$-structure $\mathfrak{A}$ through an equivalence relation $E\subseteq A^{2}$ is the $\tau$-structure $\mathfrak{A}/{E}$ with domain $A/E$ and relations $R^{\mathfrak{A}/{E}}=q_{E}(R^{\mathfrak{A}})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$.
Structure-preserving maps.
@@
A homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ for $\tau$-structures $\mathfrak{A}$ and $\mathfrak{B}$ is a mapping $h\colon A\rightarrow B$ that preserves each $\tau$-relation, i.e., if $\bar{t}\in R^{\mathfrak{A}}$ for some relation symbol $R\in\tau$, then $h(\bar{t})\in R^{\mathfrak{B}}$. We write $\mathfrak{A}\rightarrow\mathfrak{B}$ if $\mathfrak{A}$ maps homomorphically to $\mathfrak{B}$; formally, $\mathrm{CSP}(\mathfrak{A})=\{\mathfrak{J}\text{ finite}\mid\mathfrak{J}\rightarrow\mathfrak{A}\}$. If $\mathfrak{A}\rightarrow\mathfrak{B}$ and $\mathfrak{B}\rightarrow\mathfrak{A}$, then we say that $\mathfrak{A}$ and $\mathfrak{B}$ are homomorphically equivalent. An endomorphism of $\mathfrak{A}$ is a homomorphism from $\mathfrak{A}$ to itself; we denote by $\mathrm{End}(\mathfrak{A})$ the set (monoid) of all endomorphisms of $\mathfrak{A}$. It is easy to see that if $\mathfrak{A}$ has a constant endomorphism, then its CSP is trivial, and we call $\mathfrak{A}$ trivial itself; otherwise we call $\mathfrak{A}$ non-trivial.

### v2

@@ 2.2. Polymorphisms are injective
The CSP of $(\mathbb{Q};<)_{I_{4}}$ from Example 2.1 is solvable in Datalog (Section 4.1). Since $\mathrm{CSP}(\mathfrak{K}_{3})$ is not, and pp-constructions provide Datalog-reductions [1], it follows that $(\mathbb{Q};<)_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. Hence its polymorphisms satisfy the pseudo-Siggers identity by the above-mentioned theorem of Barto and Pinsker. However, any polymorphism of $I_{4}$ is injective up to dummy variables [16], and hence cannot satisfy the Siggers identity.
We say that a template is Pol-injective if all of its polymorphisms are essentially injective, i.e. injective up to dummy variables – this is the case if and only if the relation $I_{4}$ is invariant under its polymorphisms. As it turns out, the relation $I_{4}$ can be added to the templates constructed in Theorem 2.2 at no computational cost (up to Datalog reductions). It thus follows that any algebraic condition on polymorphisms capturing a complexity class closed under Datalog-reductions for CSPs of structures within the scope of Conjecture 1.4 must necessarily be satisfiable by essentially injective operations. Until now, this fact was understood by experts on an intuitive level, but our result provides a rigorous confirmation.
@@
The structures $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}}$ enjoy all properties of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ from Theorem 2.2 except that CSP-injectivity is replaced by Pol-injectivity.
For finite structures the only essentially injective polymorphisms are at most unary up to dummy variables. It is therefore of no surprise that so few universal-algebraic conditions for CSPs from the finite have been successfully lifted to the infinite. One prominent example which can actually be lifted is the satisfiability of quasi near-unanimity identities, which captures bounded strict width both in the finite and in the infinite [17]; bounded strict width corresponds to solvability in a particular proper fragment of Datalog that is not closed under the Datalog-reductions in Theorem 2.2. Examples of conditions for polymorphism clones that can be satisfied by essentially injective operations are any pseudo height-1 identities (e.g., the pseudo-Siggers identity), but also the dissected near-unanimity identities [47] which were proposed in [33] as a possible candidate for solvability of CSPs in fixed-point logic.
Marimon and Pinsker have recently provided a negative answer, within the scope of Conjecture 1.4, to a question of Bodirsky [15, Question 14.2.6(27)] asking whether every $\omega$-categorical CSP template without algebraicity which is solvable in Datalog has a binary injective polymorphism [55, Corollary 8.10]. We contrast this negative result with a corollary to Theorem 2.4 that implies that up to Datalog reductions the answer is positive.
@@
Let $\mathfrak{A}$ be a non-trivial reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. If $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, then $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has a binary injective polymorphism.
2.3. Everything is a cheese
@@
One of the currently most vibrant branches of research in constraint satisfaction are Promise Constraint Satisfaction Problems (PCSPs). For two relational structures $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ such that $\mathfrak{S}_{1}$ maps homomorphically to $\mathfrak{S}_{2}$, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to decide whether a given finite structure $\mathfrak{J}$ in the same signature as $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ maps homomorphically to $\mathfrak{S}_{1}$ or does not even map homomorphically to $\mathfrak{S}_{2}$; the instance $\mathfrak{J}$ is promised to satisfy one of those two cases. Thus, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to compare $\mathrm{CSP}(\mathfrak{S}_{1})$ to a relaxation $\mathrm{CSP}(\mathfrak{S}_{2})$ of it, and since $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{1})=\mathrm{CSP}(\mathfrak{S}_{1})$, promise constraint satisfaction problems generalize the CSP framework. This generalization is vast, and contains many well-known computational problems such as APPROXIMATE GRAPH COLORING [45], $(2+\varepsilon)$-SAT [2], and HYPERGRAPH COLORING [40]. Although formally not necessary, research has focused on finite templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$: this is justified for example by the fact that even for structures over a Boolean domain, there is no complete complexity classification yet. Hence, the two extensions considered here of finite-domain CSPs to $\omega$-categorical templates, or alternatively to PCSPs, are somewhat orthogonal, and it is natural to wonder whether there exist any connections, in particular of algorithmic nature.
@@
One potential link is the use of $\omega$-categorical CSP templates as cheeses in the sandwich method to prove tractability of PCSPs, as follows. Note that if $\mathfrak{A}$ is a structure such that $\mathfrak{S}_{1}$ maps homomorphically to $\mathfrak{A}$ and $\mathfrak{A}$ maps homomorphically to $\mathfrak{S}_{2}$, then $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ trivially reduces to $\mathrm{CSP}(\mathfrak{A})$; we call $\mathfrak{A}$ a cheese sandwiched by the template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$. This simple observation is more powerful than it might first appear: in fact, so far, every known tractable (finite-domain) PCSP can be reduced to the tractable CSP of some cheese structure via the sandwich method [56, 39]. On the other hand, Barto [4] provided an example of a PCSP template sandwiching an infinite polynomial-time tractable cheese while not being finitely tractable, i.e. such that every finite-domain cheese has an NP-complete CSP [4]. Since Barto’s cheese is not $\omega$-categorical, he raised the question whether such an example could also be constructed with an $\omega$-categorical cheese ([4, Question IV.1]). Subsequently the question arose, and was promoted in particular by Zhuk at the CSP World Congress 2023, whether structures within the scope of Conjecture 1.4 could ever serve as tractable cheeses in the absence of a finite one.
@@
The first two examples of tractable $\omega$-categorical cheeses for non-finitely tractable finite-domain PCSPs were recently obtained by Mottet [57]; his structures fall within the scope of Conjecture 1.4. In the present paper, we strengthen this result by creating non-finitely tractable finite-domain PCSPs from virtually every structure $\mathfrak{A}$ within the scope of Conjecture 1.4. Our construction consists of two steps. First, we show that such a construction is possible in general under the additional assumption that $\mathfrak{A}$ is equipped with an inequality predicate and that it is a reduct of a structure $\mathfrak{B}$ which is linearly ordered. We then combine this construction with Theorem 2.2. Modulo an extra step consisting of taking generic superpositions (Proposition 5.3) of structures with $(\mathbb{Q};<)$ (for which we need the property of no algebraicity), this finishes our proof.
Theorem 2.6 (Datalog-Approximability of Infinite-Domain CSPs).
@@
For every non-trivial $\tau$-reduct $\mathfrak{A}$ of a countable finitely bounded homogeneous structure $\mathfrak{B}$ there exists a $\tau^{\prime}$-reduct $\mathfrak{A}^{\prime}$ of a countable finitely bounded homogeneous structure $\mathfrak{B}^{\prime}$ and Datalog-interpretations $\mathcal{J},\mathcal{J}^{\prime}$ from $\tau$-structures to $\tau^{\prime}$-structures and vice versa such that, for every $n\in\mathbb{N}$, there exists a finite PCSP template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ with the following properties:
@@
    $\mathcal{J}$ and $\mathcal{J}^{\prime}$ are (Datalog-)reductions from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}(\mathfrak{A}^{\prime})$ and vice versa.
@@
  5. (5)
    Every finite cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ pp-constructs $\mathfrak{K}_{3}$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is not finitely tractable.
  6. (6)
    $\mathfrak{A}^{\prime}$ is a cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ reduces to $\mathrm{CSP}(\mathfrak{A})$ via $\mathcal{J}^{\prime}$.
  7. (7)
    Conversely, $\mathcal{J}$ is correct as a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ when limited to instances $\mathfrak{X}$ of $\mathrm{CSP}(\mathfrak{A})$ with the property that $\mathfrak{X}$ homomorphically maps to $\mathfrak{A}$ if and only if $\mathfrak{X}$ homomorphically maps to a substructure of $\mathfrak{A}$ of domain size $\leq n$.
The last item in Theorem 2.6 is not necessary for obtaining non-finitely tractable PCSPs from Conjecture 1.4, but provides an additional link that makes it possible to transfer logical inexpressiblity results from infinite-domain CSPs to PCSPs. For example, the PCSPs provided by the theorem from $\mathrm{CSP}(\mathbb{Q};X)$, where
$X=\{(x,y,z)\mid x=y<z\text{ or }y=z<x\text{ or }z=x<y\}$,
will not be solvable in Fixed-Point logic with Counting (FPC), because $\mathrm{CSP}(\mathbb{Q};X)$ is not [33] and the instances witnessing the inexpressiblity in FPC homomorphically map to $(\mathbb{Q};X)$ if and only if they homomorphically map to a fixed finite substructure of it (cf. [57, Proposition 37]). In contrast, $\mathrm{CSP}(\mathbb{Q};\{(x,y,z)\mid x>y\text{ or }x>z\})$ is not solvable in Datalog [23], but it can be deduced from the results in [57] (Theorem 3) that the associated finite-domain PCSPs are.
@@
As mentioned earlier, the first two examples of a tractable $\omega$-categorical cheese for a finite-domain PCSP that is not finitely tractable have recently been provided by Mottet [57, Propositions 35 and 36]. Our results compare to the results of that work as follows. First, Mottet’s finite-domain PCSPs [57, Theorem 2] are reducible to the associated $\omega$-categorical templates under gadget reductions, which are the specific form of Datalog-reductions stemming from pp-constructions. The main advantage of gadget reductions over general Datalog-reductions is that, in many settings, gadget reductions between (P)CSPs are captured by the transfer of height-1 identities between polymorphism clones (or more generally polymorphism minions); see [38]. As a trade-off for the generality of our result (Theorem 2.6), our Datalog-reductions are not given in terms of gadgets; in fact, one can formally prove that, in many cases, our finite-domain PCSPs do not admit a gadget reduction to the original $\omega$-categorical template (Example 4.6). Second, Mottet’s work [57] also contains a result (Theorem 1) showing that every CSP within the scope of the Bodirsky-Pinsker conjecture is polynomial-time equivalent to the PCSP of a half-infinite template, more precisely a template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that $\mathfrak{S}_{1}$ is in the scope of the Bodirsky-Pinsker conjecture and $\mathfrak{S}_{2}$ is finite. In contrast, our Theorem 2.6 provides non-finitely tractable finite-domain PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ sandwiching a given structure $\mathfrak{A}$ from Conjecture 1.4 up to Datalog interreducibility. Since our PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ are finite and not half-infinite, we cannot guarantee polynomial-time equivalence between $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ and $\mathrm{CSP}(\mathfrak{A})$. However, our construction still allows us to approximate $\mathrm{CSP}(\mathfrak{A})$ to an arbitrary degree in the sense of item 7 in Theorem 2.6.
3. Basic definitions
@@
The set $\{1,\dots,n\}$ is denoted by $[n]$, and we use the bar notation $\bar{t}$ for tuples. For a function $f\colon A^{n}\rightarrow B$, the component-wise action of $f$ on $k$-tuples $(x_{1,1},\dots,x_{1,k}),$ $\dots$,… , $(x_{n,1},\dots,x_{n,k})\in A^{k}$ is defined as the $k$-tuple
$f\big{(}(x_{1,1},\dots,x_{1,k}),\dots,(x_{n,1},\dots,x_{n,k})\big{)}\coloneqq\big{(}f(x_{1,1},\dots,x_{n,1}),\dots,f(x_{1,k},\dots,x_{n,k})\big{)}.$
3.1. Relational structures.
@@
A (relational) signature $\tau$ is a set of relation symbols, each $R\in\tau$ with an associated natural number called arity. A (relational) $\tau$-structure $\mathfrak{A}$ consists of a set $A$ (the domain) together with the relations $R^{\mathfrak{A}}\subseteq A^{k}$ for each $R\in\tau$ with arity $k$. An expansion of $\mathfrak{A}$ is a $\sigma$-structure $\mathfrak{B}$ with $A=B$ such that $\tau\subseteq\sigma$ and $R^{\mathfrak{B}}=R^{\mathfrak{A}}$ for each relation symbol $R\in\tau$. Conversely, we call $\mathfrak{A}$ a reduct of $\mathfrak{B}$ and denote it by $\mathfrak{B}|^{\tau}$. For a positive integer $d$ and a $\tau$-structure $\mathfrak{A}$, the $d$-th power of $\mathfrak{A}$ is the $\tau$-structure $\mathfrak{A}^{d}$ with domain $A^{d}$ and relations
$R^{\mathfrak{A}^{d}}=\{\big{(}(a_{1,1},\dots,a_{1,d}),\dots,(a_{k,1},\dots,a_{k,d})\big{)}\in(A^{d})^{k}\mid(a_{1,1},\dots,a_{k,1}),\dots,(a_{1,d},\dots,a_{k,d})\in R^{\mathfrak{A}}\}$
for each $R\in\tau$ of arity $k$. The substructure of a $\tau$-structure $\mathfrak{A}$ on a subset $B\subseteq A$ is the $\tau$-structure $\mathfrak{B}$ with domain $B$ and relations $R^{\mathfrak{B}}=R^{\mathfrak{A}}\cap B^{k}$ for every $R\in\tau$ of arity $k$. The factor of a $\tau$-structure $\mathfrak{A}$ through an equivalence relation $E\subseteq A^{2}$ is the $\tau$-structure $\mathfrak{A}/{E}$ with domain $A/E$ and relations $R^{\mathfrak{A}/{E}}=q_{E}(R^{\mathfrak{A}})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$.
@@
A homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ for $\tau$-structures $\mathfrak{A}$ and $\mathfrak{B}$ is a mapping $h\colon A\rightarrow B$ that preserves each $\tau$-relation, i.e., if $\bar{t}\in R^{\mathfrak{A}}$ for some relation symbol $R\in\tau$, then $h(\bar{t})\in R^{\mathfrak{B}}$. We write $\mathfrak{A}\rightarrow\mathfrak{B}$ if $\mathfrak{A}$ maps homomorphically to $\mathfrak{B}$; formally, $\mathrm{CSP}(\mathfrak{A})=\{\mathfrak{J}\text{ finite}\mid\mathfrak{J}\rightarrow\mathfrak{A}\}$. If $\mathfrak{A}\rightarrow\mathfrak{B}$ and $\mathfrak{B}\rightarrow\mathfrak{A}$, then we say that $\mathfrak{A}$ and $\mathfrak{B}$ are homomorphically equivalent. An endomorphism of $\mathfrak{A}$ is a homomorphism from $\mathfrak{A}$ to itself; we denote by $\mathrm{End}(\mathfrak{A})$ the set (monoid) of all endomorphisms of $\mathfrak{A}$. Clearly, if $\mathfrak{A}$ has a constant endomorphism, then its CSP is trivial, and we call $\mathfrak{A}$ trivial itself; otherwise we call $\mathfrak{A}$ non-trivial. As mentioned in the introduction, a polymorphism of $\mathfrak{A}$ is a homomorphism from $\mathfrak{A}^{k}$ into $\mathfrak{A}$ for some $k\in\mathbb{N}$.